OpenStudy (anonymous):

I have no idea how to do this problem. Any help would be greatly appreciated. Consider the function f(x) for which f(0)=4 and f'(0)=11. Find h'(0) for the function h(x)=1/(f(x)

hartnn (hartnn):

whats the function h(x) again ??

hartnn (hartnn):

anyways, you'll need chain rule here, you know whats a chain rule ?

OpenStudy (anonymous):

To be honest I know the formula but I don't know how to use is properly. My professor was gone the last 4 days and the person we had teaching was foreign and I couldn't understand him.

hartnn (hartnn):

ok, tell me again what h(x) is ?

OpenStudy (anonymous):

h(x)=1/f(x)

hartnn (hartnn):

ok, tell me deribative of 1/x first ?

OpenStudy (anonymous):

-1/x^2

hartnn (hartnn):

ok, so using chain rule, derivative of 1/ f(x) will be -1/f(x)^2 * [f'(x)] got this ?

OpenStudy (anonymous):

So how do I use that to find h'(0)

hartnn (hartnn):

that was h'(x) =-1/f(x)^2 * [f'(x)] now just put x=0 and you have both, f(0) ans f'(0)

OpenStudy (anonymous):

Oh I understand. Thank you very much for your help

hartnn (hartnn):

what you got as answer ? just to verify...

OpenStudy (anonymous):

-11/16

hartnn (hartnn):

thats correct :) welcome ^_^

OpenStudy (anonymous):

you sir saved my life. Thanks again

hartnn (hartnn):

your are welcomed anytime :)